# A Comprehensive Survey of Machine Unlearning Techniques for Large Language Models

### v1

@@ Abstract
This study investigates the machine unlearning techniques within the context of large language models (LLMs), referred to as LLM unlearning. LLM unlearning offers a principled approach to removing the influence of undesirable data (e.g., sensitive or illegal information) from LLMs, while preserving their overall utility without requiring full retraining. Despite growing research interest, there is no comprehensive survey that systematically organizes existing work and distills key insights; here, we aim to bridge this gap. We begin by introducing the definition and the paradigms of LLM unlearning, followed by a comprehensive taxonomy of existing unlearning studies. Next, we categorize current unlearning approaches, summarizing their strengths and limitations. Additionally, we review evaluation metrics and benchmarks, providing a structured overview of current assessment methodologies. Finally, we outline promising directions for future research, highlighting key challenges and opportunities in the field.
@@ 1 Introduction
Figure 2: Taxonomy of LLM unlearning.
The widespread adoption of large language models (LLMs) has brought significant challenges, particularly concerning user data privacy, copyright protection, and alignment with societal values. During training, these models can inadvertently memorize sensitive information, such as personally identifiable data or copyrighted materials Li et al. (2024a, b); Zhang et al. (2024b); Yao et al. (2024). In addition to privacy and copyright issues, some training data may embed content that conflicts with contemporary social norms, such as discriminatory language based on race, ethnicity, etc Li et al. (2025). These biases often manifest as harmful stereotypes, undermining the fairness and inclusivity of AI systems. Addressing these concerns is not only a societal imperative but also a regulatory requirement under privacy laws such as the General Data Protection Regulation (GDPR) Council of the European Union (2016) and the EU Artificial Intelligence Act (EU AI Act) Council of the European Union (2024). These laws mandate the “right to be forgotten” and require mechanisms to delete specific data upon request.
To address these challenges, the field of LLM unlearning has emerged, focusing on removing specific information or behaviors from models while preserving their overall performance. However, LLM unlearning faces significant technical challenges. One of the most pressing issues is the prohibitively high cost of retraining. Traditionally, addressing harmful or unwanted data required retraining the model from scratch after excluding problematic data from the training set Jang et al. (2023). This is impractical for LLMs due to the immense time and computational resources required Li et al. (2024a). Moreover, the frequent unlearning requests that arise in deployed models highlight the need for more efficient unlearning techniques. The complexity of LLMs, with their millions or even hundreds of billions parameters, further complicates the task of removing specific information without causing unintended side effects, such as performance degradation or catastrophic forgetting Zhang et al. (2024a).
Despite recent advances in LLM unlearning Eldan et al. (2024); Maini et al. (2024), there is no comprehensive review of existing approaches and evaluation measures. Here we aim to bridge this gap. In particular, we offer a thorough overview of the field, including various unlearning and evaluation methods, and we make the following contributions:
@@
  • We systematically review the existing methods, analyzing their features, strengths, and weaknesses. We further discuss the existing evaluation measures and benchmarks, highlighting the challenges when balancing utility preservation and forgetting quality.
  • We discuss future research opportunities for LLM unlearning, including extending techniques to multimodal models and addressing real-world unlearning requests. These avenues aim to advance the field and address emerging challenges.
@@ 2.2 LLM Unlearning Paradigms
LLM unlearning follows two main paradigms. The fine-tuning-then-unlearning paradigm focuses on eliminating knowledge introduced during fine-tuning. As illustrated in Figure 1, this paradigm typically leverages synthetic data (e.g., TOFU Maini et al. (2024) and FIUBench Ma et al. (2024)) and partitions a task-specific dataset into a forget set $\mathcal{D}_{f}$ and a retain set $\mathcal{D}_{r}$ to highlight the unlearning precision. The original model $\theta_{ori}$ is first obtained by fine-tuning a pretrained model $\theta_{pre}$ on the task-specific data to encode the target knowledge. Unlearning techniques are then applied to reduce the model’s reliance on the forget set, resulting in the unlearned model $\theta_{unl}$. The direct-unlearning paradigm focuses on eliminating knowledge acquired during the pretraining stage of $\theta_{ori}$, assuming the target knowledge originates from multiple points within the pretraining dataset. The forget and retain sets are typically sampled from pretraining corpora Yao et al. (2024) or publicly available datasets such as Wiki Jin et al. (2024). This paradigm is also extensively applied in safety alignment tasks, where it aims to eradicate hazardous knowledge and to mitigate risks such as misuse or jailbreak attacks Li et al. (2024a); Zhang et al. (2024b).
Substitutes for Knowledge to be Forgotten
Existing unlearning paradigms primarily specify the knowledge to be forgotten, while leaving the substitutes undefined. In this case, refusal responses, such as “I don’t know the answer,” are commonly used as substitutes Maini et al. (2024); Ma et al. (2024)Yao et al. [2024] explored two relabeling strategies. The first strategy uses random labels generated by a randomly initialized model, while the second involves sampling adversarial examples from the model. For each token $w_{t}$ to be unlearned, an adversarial sample $a_{t}$ is generated to maximize the model’s confusion while staying close to $w_{t}$. Eldan et al. [2024] proposed a fine-grained relabeling approach. Their method first identifies tokens whose logit probabilities increase after fine-tuning on the forget set. To facilitate unlearning, a dictionary generated by GPT-4 replaces specific entities with contextually appropriate generic alternatives (e.g., replacing “Hogwarts” with “Mystic Academy”), ensuring more natural and context-aware substitutions.
@@ 2.3 Taxonomy
We present a comprehensive taxonomy of LLM unlearning, as illustrated in Figure 2, outlining existing research from the perspectives of methods, evaluation measures, and benchmarks. Existing methods can be categorized into four types: direct fine-tuning, localized parameter modification, leveraging auxiliary models, and input/output-based unlearning. Forgetting quality and utility preservation are critical measures for evaluating unlearning algorithms, particularly given recent discussions on whether knowledge is robustly forgotten or remains susceptible to adversarial recovery. This is often assessed through input-based or logit-based evaluation, as well as model intervention techniques. Additionally, we review commonly used unimodal and multimodal benchmarks.
@@ 3.1 Direct Fine-Tuning
Relabeling-Based Fine-Tuning first replaces the original responses by generic or neutral substitutes, as described earlier. Then, the model is fine-tuned on the relabeled data to reduce the influence of undesired training information Jin et al. (2024); Eldan et al. (2024).
Reinforcement Learning
@@
Gradient Ascent (GA)
Jang et al. [2023] performed gradient ascent (GA) on the next-token prediction loss over the forget set, which is mathematically equivalent to performing gradient descent on the negative prediction loss:
$\small\mathcal{L}_{GA}(\theta)=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\left[-\log\left(p(y\mid x;\theta)\right)\right].$ (2)
However, the GA-based approach often results in sizable model degradation, characterized by the generation of uniform and low-quality outputs. To mitigate this and to preserve model utility,  Liu et al. [2022] introduced the gradient descent (GD) loss on the retain set $\mathcal{D}_{r}$ as a regularization:
$\small\mathcal{L}_{GD}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{r}}\left[-\log\left(p(y\mid x;\theta)\right)\right].$ (3)
Yao et al. [2024] incorporated the Kullback-Leibler (KL) divergence term as the regularization, which captures the difference between the predictions of the original model and the fine-tuned model on the retain set:
$\small\textstyle\mathcal{L}_{KL}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{r}}\left[D_{KL}\left(p(y\mid x;\theta_{ori})\,\|\,p(y\mid x;\theta)\right)\right].$ (4)
These regularization techniques can also be applied in other unlearning methods to preserve the model’s utility.
Preference Optimization was first designed to align model behavior to human-defined preferences. Specifically, it leverages pairwise comparisons or ranking data to guide the model toward producing outputs that best match desired preferences. Given a preference dataset $\mathcal{D}_{p}=\left\{(x_{i},y_{i,w},y_{i,l})\right\}_{i\in[n]}$, where $y_{i,w}$ and $y_{i,l}$ represent responses to input $x_{i}$, the preference $y_{i,w}>y_{i,l}$ is derived from human comparisons. Direct preference optimization (DPO) Rafailov et al. (2023) minimizes the following objective function:
@@
$\small\textstyle\mathcal{L}_{\text{DPO},\beta}(\theta)=-\frac{2}{\beta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\left[\log\sigma\left(-\beta\log\frac{p(y\mid x;\theta)}{p(y\mid x;\theta_{ori})}\right)\right].$ (6)
@@ 3.2 Localized Parameter Modification
RMU Li et al. (2024a) focuses on unlearning hazardous knowledge in LLMs by fine-tuning the lower layer $l$ to redirect internal representations of token $t$ in the forget set toward a fixed-noise vector $\mathbf{u}$:
$\small\mathcal{L}_{\text{forget}}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\frac{1}{L_{x}}\sum_{t\in x}\|\mathcal{H}^{(l)}_{\theta}(t)-c\cdot\mathbf{u}\|_{2}^{2},$ (7)
where $L_{x}$ is the number of tokens in $x$ and $c$ is some hyperparameter that controls activation scaling, $\mathcal{H}^{(l)}_{\theta}(t)$ denotes the internal activations of token $t$ at layer $l$. Simultaneously, it ensures that preserved knowledge remains consistent with the original model by aligning its representations, denoted as
$\small\mathcal{L}_{\text{retain}}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{r}}\frac{1}{L_{x}}\sum_{t\in x}\|\mathcal{H}^{(l)}_{\theta}(t)-\mathcal{H}^{(l)}_{\theta_{\text{ori}}}(t)\|_{2}^{2}.$ (8)
Huu-Tien et al. [2024] proposed adaptive RMU, which dynamically scales the random unit vector $\mathbf{u}$ with an adaptive coefficient $\beta\|\mathcal{H}_{\theta_{\text{ori}}}^{(l)}(x)\|_{2}^{2}$ for improved unlearning across layers, unlike RMU’s fixed-scaling coefficient $c$.
@@
WAGLE Jia et al. (2024) uses bi-level optimization to examine the interaction between weight adjustment and unlearning efficacy. By leveraging weight attribution, it quantifies the relationship between weight influence and the impact of forgotten or retained data on LLM outputs. The unlearning sensitivity score for weight perturbation is obtained from the forget loss $S_{i}=\mathcal{L}_{f}(\epsilon\odot\theta(\epsilon))-\mathcal{L}_{f}(\theta(1))$, where $\epsilon\odot\theta(\epsilon)$ is the weight-adjusted model, $\epsilon$ represents weight modifications, and $\epsilon=1$ indicates no interference. The weights $\theta(\epsilon)$ minimize the retain loss $\mathcal{L}_{r}$ to preserve utility. WAGLE uses a diagonal Hessian approximation for computational efficiency and accuracy, with the sensitivity score expressed as
@@
Table 1: A summary of existing LLM unlearning studies. C-Utility, C-Forget, C-Efficiency denote the challenges to solve in the perspective of utility preservation, forget quality, and efficiency, respectively.
Methodology Features C-Utility C-Forget C-Efficiency Related Work Application
Relabeling-Based Fine-Tuning Using refusal string or fine-grained alternatives Eldan et al. (2024) Privacy & Copyright
Gradient Ascent Maximizing cross-entropy loss w.r.t target sequence Jang et al. (2023) Privacy & Copyright
Yao et al. (2024) Privacy & Copyright
Preference Optimization Better utility preservation Maini et al. (2024) Privacy & Copyright
Zhang et al. (2024a) Privacy & Copyright
Zhang et al. (2024b) AI Alignment
Reinforcement Learning Need for an accurate reward model Kassem et al. (2023) Privacy & Copyright
Lu et al. (2022) AI Alignment
Representation Engineering Modify localized weights to update internal representation Li et al. (2024a) AI Alignment
Huu-Tien et al. (2024) AI Alignment
Locate-then- Unlearn First locate the parameters corresponding to the target knowledge, then modify them. Wu et al. (2023) Privacy & Copyright
Jia et al. (2024) Privacy & Copyright
Guo et al. (2024) Privacy & Copyright
Hong et al. (2024) Privacy & Copyright
Task Vector Utilize the model weight difference before and after fine-tuning Ilharco et al. (2023) AI Alignment
Liu et al. (2024c) AI Alignment
Gao et al. (2024) AI Alignment
Contrastive Decoding Intervention during the decoding process Ji et al. (2024) AI Alignment
Knowledge Distillation Adjusting logits to guide the student model Wang et al. (2024) Privacy & Copyright
Dong et al. (2024) Privacy & Copyright
Input/Output- Based Modify inputs or outputs instead of model weights. Pawelczyk et al. (2025) Privacy & Copyright
Liu et al. (2024a) Privacy & Copyright
Thaker et al. (2024) Privacy & Copyright

### v2

# A Survey of Machine Unlearning in Large Language Models: Methods, Challenges and Future Directions

@@ Abstract
This study investigates the machine unlearning techniques within the context of large language models (LLMs), referred to as LLM unlearning. LLM unlearning offers a principled approach to removing the influence of undesirable data (e.g., sensitive or illegal information) from LLMs, while preserving their overall utility without requiring full retraining. Despite growing research interest, there is no comprehensive survey that systematically organizes existing work and distills key insights; here, we aim to bridge this gap. We begin by introducing the definition and the paradigms of LLM unlearning, followed by a comprehensive taxonomy of existing unlearning studies. Next, we categorize current unlearning approaches, summarizing their strengths and limitations. Additionally, we review evaluation measures and benchmarks, providing a structured overview of current assessment methodologies. Finally, we outline promising directions for future research, highlighting key challenges and opportunities in the field.
A Survey of Machine Unlearning in Large Language Models:
Methods, Challenges and Future Directions
Qing Li1†  Jiahui Geng1†  Herbert Woisetschläger3  Zongxiong Chen2  Fengyu Cai4 Yuxia Wang1Preslav Nakov1Hans Arno Jacobsen5Fakhri Karray1 1Mohamed bin Zayed University of Artificial Intelligence (MBZUAI), UAE 2Fraunhofer Institute for Open Communication Systems (FOKUS), Germany 3Technical University of Munich, Germany 4Technical University of Darmstadt, Germany 5University of Toronto, Canada
@@ 1 Introduction
Related Work Date Taxonomy # Methodologies Modality Adv Evaluation
Si et al. (2023) December 2023 Yes 3 Unimodal 0
Xu (2024) April 2024 Yes 3 Unimodal 0
Liu et al. (2025) June 2024 No 6 Unimodal,Multimodal 5
Liu et al. (2024c) July 2024 No 5 Unimodal,Multimodal 0
Ours May 2025 Yes 10 Unimodal,Multimodal 7
Table 1: Comparison of different surveys on the LLM Unlearning. Adv Evaluation refers to Adversarial Evaluation.
[…]
mitigate_leaf/.style= fill=green!10, draw=green!30, rounded corners=2pt, text=black , detect_leaf/.style= fill=blue!10, draw=blue!30, rounded corners=2pt, text=black , hidden-draw/.style= draw=none ,
The widespread adoption of large language models (LLMs) has brought significant challenges, particularly concerning user data privacy, copyright protection, and alignment with societal values. During training, these models can inadvertently memorize sensitive information, such as personally identifiable data or copyrighted materials Li et al. (2024b, c); Zhang et al. (2024b); Yao et al. (2024). In addition to privacy and copyright issues, some training data may embed content that conflicts with contemporary social norms, such as discriminatory language based on race, ethnicity, etc Li et al. (2025a). These biases often manifest as harmful stereotypes, undermining the fairness and inclusivity of AI systems. Addressing these concerns is not only a societal imperative but also a regulatory requirement under privacy laws such as the General Data Protection Regulation (GDPR, Council of the European Union 2016) and the EU Artificial Intelligence Act (EU AI Act, Council of the European Union 2024). These laws mandate the “right to be forgotten" and require mechanisms to delete specific data upon request.
To address these challenges, the field of LLM unlearning has emerged, focusing on removing specific information or behaviors from models while preserving their overall performance. However, LLM unlearning faces significant technical challenges. One of the most pressing issues is the prohibitively high cost of retraining. Traditionally, addressing harmful or unwanted data required retraining the model from scratch after excluding problematic data from the training set Jang et al. (2023). This is impractical for LLMs due to the immense time and computational resources required Li et al. (2024b). Moreover, the frequent unlearning requests that arise in deployed models highlight the need for more efficient unlearning techniques. The complexity of LLMs, with their millions, or even hundreds of billions of parameters, further complicates the task of removing specific information without causing unintended side effects, such as performance degradation or catastrophic forgetting Zhang et al. (2024a).
Table 1 provides a comparative overview of existing survey papers on LLM unlearning alongside our work. Most prior surveys were published before July 2024 and therefore do not capture recent advancements in the field. Moreover, these surveys either lack a systematic taxonomy or are limited to unimodal unlearning methods. Here we aim to bridge this gap. In particular, we offer a thorough overview of the field, including various unlearning and evaluation methods, and we make the following contribution
@@
  • We systematically review the existing methods, analyzing their strengths and weaknesses. We further examine the existing evaluation measures and benchmarks, highlighting the challenges of balancing utility preservation with forgetting quality.
  • We discuss future research opportunities for LLM unlearning, including extending techniques to multimodal models and addressing complex real-world unlearning requests. These avenues aim to advance the field and address emerging challenges.
@@ 2.2 LLM Unlearning Paradigms
LLM unlearning follows two main paradigms. The fine-tuning-then-unlearning paradigm focuses on eliminating knowledge introduced during fine-tuning. As illustrated in Figure 1, this paradigm typically leverages synthetic data (e.g., TOFU Maini et al. (2024) and FIUBench Ma et al. (2024)) and partitions a task-specific dataset into a forget set $\mathcal{D}_{f}$ and a retain set $\mathcal{D}_{r}$ to highlight the unlearning precision. The original model $\theta_{ori}$ is first obtained by fine-tuning a pretrained model $\theta_{pre}$ on the task-specific data to encode the target knowledge. Unlearning techniques are then applied to reduce the model’s reliance on the forget set, resulting in the unlearned model $\theta_{unl}$. The direct-unlearning paradigm focuses on eliminating knowledge acquired during the pretraining stage of $\theta_{ori}$, assuming the target knowledge originates from multiple points within the pretraining dataset. The forget and retain sets are typically sampled from pretraining corpora Yao et al. (2024) or publicly available datasets such as Wiki Jin et al. (2024). This paradigm is also extensively applied in safety alignment tasks, where it aims to eradicate hazardous knowledge and to mitigate risks such as misuse or jailbreak attacks Li et al. (2024b); Zhang et al. (2024b).
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ Machine Unlearning in LLMs, ver, color=white, fill=rootcolor, text=white, font=[ Methods, color=white, fill=methodcolor, text width=5.0em, text=white, font=[ Direct Fine-Tuning, color=black, fill=methodcolor!80, text width=7em, text=black [ Relabeling-based Fine-Tuning, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Eldan et al. (2024); Li et al. (2024a), detect_leaf, text width=20em, fill=methodcolor!20 ] ] [ Gradient Ascent, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Jang et al. (2023); Yao et al. (2024), detect_leaf, text width=20em, fill=methodcolor!20 ] ] [ Preference Optimization, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Maini et al. (2024); Zhang et al. (2024a, b), detect_leaf, text width=20em, fill=methodcolor!20 ] ] [ Reinforcement Learning, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Kassem et al. (2023); Lu et al. (2022), detect_leaf, text width=20em, fill=methodcolor!20 ] ] ] [ Localized Parameter
Modification, color=black, fill=methodcolor!80, text width=7em, text=black [ Representation Engineering, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Li et al. (2024b); Dang et al. (2025); Li et al. (2025b), detect_leaf, text width=20em, fill=methodcolor!20 ] ] [ Locate-then-Unlearn, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Wu et al. (2023); Jia et al. (2024); Hong et al. (2024), detect_leaf, text width=20em, fill=methodcolor!20 ] ] ] [ Leveraging Auxiliary
Models, color=black, fill=methodcolor!80, text width=7em, text=black [ Task Vector, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Ilharco et al. (2023); Lu et al. (2024); Gao et al. (2024), detect_leaf, text width=20em, fill=methodcolor!20 ] ] [ Contrastive Decoding, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Ji et al. (2024), detect_leaf, text width=20em, fill=methodcolor!20 ] ] [ Knowledge Distillation, color=black, fill=methodcolor!60, text width=10.0em, text=black [ Wang et al. (2024); Dong et al. (2024), detect_leaf, text width=20em, fill=methodcolor!20 ] ] ] [ Input/Output-Based
Unlearning, color=black, fill=methodcolor!80, text width=7em, text=black [ Thaker et al. (2024); Liu et al. (2022)
Pawelczyk et al. (2025), detect_leaf, text width=20em, fill=methodcolor!20 ] ] ] [ Evaluation
Measures, color=white, fill=evalcolor, text width=5.0em, text=white, font=[ Adversarial Evaluation, color=black, fill=evalcolor!80, text width=8em, text=black [ Logit-Based Evaluation, color=black, fill=evalcolor!60, text width=8.0em, text=black [ KS-Test, color=black, fill=evalcolor!40, text width=10em, text=black [ Maini et al. (2024); Ma et al. (2024), cause_leaf, text width=11.4em, fill=evalcolor!20 ] ] [ MIA, color=black, fill=evalcolor!40, text width=10em, text=black [ Jin et al. (2024); Yao et al. (2024), cause_leaf, text width=11.4em, fill=evalcolor!20 ] ] ] [ Model Intervention, color=black, fill=evalcolor!60, text width=8.0em, text=black [ Relearning with Fine-Tuning, color=black, fill=evalcolor!40, text width=10.0em, text=black [ Hu et al. (2024); Łucki et al. (2024), cause_leaf, text width=11.4em, fill=evalcolor!20 ] ] [ LLM Quantification, color=black, fill=evalcolor!40, text width=10.0em, text=black [ Zhang et al. (2025); Łucki et al. (2024), cause_leaf, text width=11.4em, fill=evalcolor!20 ] ] [ Activation Intervention, color=black, fill=evalcolor!40, text width=10.0em, text=black [ Arditi et al. (2024); Łucki et al. (2024)
Seyitoğlu et al. (2024), cause_leaf, text width=11.4em, fill=evalcolor!20 ] ] ] [ Input-based Evaluation, color=black, fill=evalcolor!60, text width=8.0em, text=black [ Paraphrasing, color=black, fill=evalcolor!40, text width=10.0em, text=black [ Maini et al. (2024); Lynch et al. (2024), cause_leaf, text width=11.4em, fill=evalcolor!20 ] ] [ Jailbreak Prompts, color=black, fill=evalcolor!40, text width=10.0em, text=black [ Lynch et al. (2024); Łucki et al. (2024), cause_leaf, text width=11.4em, fill=evalcolor!20 ] ] ] ] [ Traditional Evaluation, color=black, fill=evalcolor!80, text width=8em, text=black [ (Probe) Accuracy/GPT-Eval/Rouge/Probability/Truth Ratio, color=black, fill=evalcolor!40, text width=25em, text=black ] ] ] [ Benchmark, color=white, fill=benchcolor, text width=5.0em, text=white, font=[ Unimodal Benchmark, color=black, fill=benchcolor!80, text width=8em, text=black [ Eldan et al. (2024); Maini et al. (2024); Li et al. (2024b)
Jin et al. (2024), mitigate_leaf, text width=20em, fill=benchcolor!20 ] ] [ Multimodal Benchmark, color=black, fill=benchcolor!80, text width=8em, text=black [ Ma et al. (2024); Liu et al. (2024b)
Dontsov et al. (2024); Li et al. (2024a), mitigate_leaf, text width=20em, fill=benchcolor!20 ] ] ] [ Application, color=white, fill=appcolor, text width=5.0em, text=white, font=[ Privacy and Copyright, color=black, fill=appcolor!80, text width=8em, text=black [ Eldan et al. (2024); Maini et al. (2024), mitigate_leaf, text width=20em, fill=appcolor!20 ] ] [ AI Alignment, color=black, fill=appcolor!80, text width=8em, text=black [ Zhang et al. (2024b); Lu et al. (2022); Li et al. (2024b), mitigate_leaf, text width=20em, fill=appcolor!20 ] ] ] ]
Figure 2: The taxonomy of machine unlearning in LLMs.
@@ 2.3 Taxonomy
We present a comprehensive taxonomy of LLM unlearning in Figure 2, outlining existing research from the perspectives of methods, evaluation measures, benchmarks, and applications. Existing methods can be categorized into four types: direct fine-tuning, localized parameter modification, leveraging auxiliary models, and input/output-based unlearning. Forgetting quality and utility preservation are critical measures for evaluating unlearning algorithms, particularly given recent discussions on whether knowledge is robustly forgotten or remains susceptible to adversarial recovery. This is often assessed through input- or logit-based evaluation, as well as model intervention techniques. Additionally, we review commonly used unimodal and multimodal benchmarks.
@@ 3.1 Direct Fine-Tuning
Relabeling-based fine-tuning first replaces the original responses with generic or neutral substitutes, as present above. The LLM is then fine-tuned on relabeled data to reduce the effect of undesired information Jin et al. (2024); Eldan et al. (2024).
Gradient ascent (GA)
Jang et al. [2023] apply gradient ascent (GA) on the next-token loss over the forget set, equivalent to minimizing the negative log-likelihood:
$\mathcal{L}_{GA}(\theta)=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\left[-\log\left(p(y\mid x;\theta)\right)\right].$ (2)
However, GA often degrades model quality, producing uniform, low-quality outputs. To mitigate this, Liu et al. [2022] add a gradient descent (GD) loss on the retain set $\mathcal{D}_{r}$ as regularization. Yao et al. [2024] instead use KL divergence to align the fine-tuned model with the original on $\mathcal{D}_{r}$:
$\mathcal{L}_{KL}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{r}}\left[D_{KL}\left(p(y\mid x;\theta_{ori})\,\|\,p(y\mid x;\theta)\right)\right].$ (3)
These regularization techniques can also be applied in other unlearning methods to preserve the utility.
Reinforcement learning (RL)
@@
Preference optimization (PO) was first designed to align model behavior to human-defined preferences. Specifically, it leverages pairwise comparisons or ranking data to guide the model toward producing outputs that best match desired preferences. Given a preference dataset $\mathcal{D}_{p}=\left\{(x_{i},y_{i,w},y_{i,l})\right\}_{i\in[n]}$, where $y_{i,w}$ and $y_{i,l}$ represent responses to input $x_{i}$, the preference $y_{i,w}>y_{i,l}$ is derived from human comparisons. Direct preference optimization (DPO) Rafailov et al. (2023) minimizes the following objective function:
@@
$\mathcal{L}_{\text{DPO},\beta}(\theta)=-\frac{2}{\beta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\left[\log\sigma\left(-\beta\log\frac{p(y\mid x;\theta)}{p(y\mid x;\theta_{ori})}\right)\right].$ (5)
@@ 3.2 Localized Parameter Modification
RMU Li et al. (2024b) focuses on unlearning hazardous knowledge in LLMs by fine-tuning the lower layer $l$ to redirect internal representations of token $t$ in the forget set toward a fixed-noise vector $\mathbf{u}$:
$\mathcal{L}_{\text{forget}}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\frac{1}{L_{x}}\sum_{t\in x}\|\mathcal{H}^{(l)}_{\theta}(t)-c\cdot\mathbf{u}\|_{2}^{2},$ (6)
where $L_{x}$ is the number of tokens in $x$ and $c$ is some hyperparameter that controls activation scaling, $\mathcal{H}^{(l)}_{\theta}(t)$ denotes the internal activations of token $t$ at layer $l$. Simultaneously, it ensures that preserved knowledge remains consistent with the original model by aligning its representations, denoted as:
$\mathcal{L}_{\text{retain}}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{r}}\frac{1}{L_{x}}\sum_{t\in x}\|\mathcal{H}^{(l)}_{\theta}(t)-\mathcal{H}^{(l)}_{\theta_{\text{ori}}}(t)\|_{2}^{2}.$ (7)
Dang et al. [2025] proposed adaptive RMU, which dynamically scales the random unit vector $\mathbf{u}$ with an adaptive coefficient $\beta\|\mathcal{H}_{\theta_{\text{ori}}}^{(l)}(x)\|_{2}^{2}$ for improved unlearning across layers, unlike RMU’s fixed-scaling coefficient $c$.
@@
WAGLE Jia et al. (2024) uses bi-level optimization to examine the interaction between weight adjustment and unlearning efficacy. By leveraging weight attribution, it quantifies the relationship between weight influence and the impact of forgotten or retained data on LLM outputs. The unlearning sensitivity score for weight perturbation is obtained from the forget loss $S_{i}=\mathcal{L}_{f}(\epsilon\odot\theta(\epsilon))-\mathcal{L}_{f}(\theta(1))$, where $\epsilon\odot\theta(\epsilon)$ is the weight-adjusted model, $\epsilon$ represents weight modifications, and $\epsilon=1$ indicates no interference. The weights $\theta(\epsilon)$ minimize the retain loss $\mathcal{L}_{r}$ to preserve utility. WAGLE uses a diagonal Hessian approximation for computational efficiency and accuracy, with the sensitivity score expressed as: